OpenStudy (anonymous):

how do I find the function when I was given the 2nd derivative?

OpenStudy (anonymous):

find the anti derivative twice example?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

integrate twice. say you're given f''(x)=2 f'(x)=2x f(x)=x^2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

well need a little more than this. \[f''(x)=2\] \[f'(x)=2x+c\] \[f(x)=x^2+cx+d\] in fact